the measure of an angle is 20 degrees less than three times its supplement. find the measure of each angle
Answer by mananth(16949)

let angle be x
supplement = 180-x
x = 3(180-x)-20
x=540-3x-20
3x+x=540-20
4x=520
x=130
supplement = 50 deg